OpenStudy (anonymous):

Simplify the following expression. √-14√-98 Here is my work did I do it right?: =√-1*√14*√-1*√98 =i*3.74*i*9.90 =1^2*37.03 =-1*37.03 = -37.03

OpenStudy (anonymous):

i think right .

OpenStudy (anonymous):

Ok, thank you. I was looking at my book and did it by it, but then let me ask you. There is a difference in √-14√-98 and between √-14*√-98 right, you do the problems differently, right?

OpenStudy (anonymous):

best thing to do is to plug it into a ti 84 and it does i's

OpenStudy (anonymous):

u right ,but u should write i^2 on second line,and subtitution i^2 =-1 in third line

OpenStudy (anonymous):

JoeJoeJame I do not have that calculator yet, I bet I will need it with my other math classes. California.Mathematics Ok, I will write it like that, thank you. Crocodile Thank you as well, and Thank you JoeJoeJame.

OpenStudy (anonymous):

√-1* √-1 = i*i = i^2 =-1